OpenStudy (alurahenderson):

What is the approximate perimeter of a semicircle with a radius of 8 m? Use pi star times = 3.14. Express your final answer to the nearest hundredth.

Vocaloid (vocaloid):

perimeter of a semi-circle = (1/2)(pi)(2)(r)

OpenStudy (alurahenderson):

4.186?

Mehek (mehek14):

\(\dfrac{1}{2}*3.14*2*8 = 25.12\)

OpenStudy (alurahenderson):

Oh I don't know how I got 4.186 I can't medal both of you so ill fan you both

OpenStudy (alurahenderson):

thank you

Mehek (mehek14):

yw